OpenStudy (abbycross167):

can someone please help me fix my math problem? please? I will fan and give a medal!! Question~ write the following inequalities in y = mx + b form. Please show all work. –3x –4y > 12 My answer~ Add 3x to both sides, -4y>12+3x Divide both sides by -4 (-4) -4y>12+3x(-4) (-4)/-4=1 3/(-4)= 0.75 = 1y>12+0.75x

OpenStudy (abbycross167):

What my teacher said was wrong~ "Watch your signs"

OpenStudy (abbycross167):

@mathstudent55 @tom982 @welshfella could one of y'll try to help me? Please?

OpenStudy (anonymous):

When you divide by a negative, you have to change the sign in the inequality. For example: -1 > -2 Divide by -1 to get 1 > 2 which is incorrect. Instead, it should be -1 > -2 goes to 1 < 2.

OpenStudy (kamibug):

In addition to Tom's comment, your final answer is not in y = mx + b form. You flipped the places of mx and b. :)

OpenStudy (kamibug):

Also, you can simplify 1y to just y.

OpenStudy (abbycross167):

So here is the only place I messed up? (-4) -4y>12+3x(-4)

OpenStudy (abbycross167):

@KamiBug

OpenStudy (kamibug):

Right, you're dividing, not multiplying. \[\frac{ -4y }{ -4 } > \frac{ 12 + 3x }{ -4 }\] This leaves you with\[y < -\frac{ 3 }{ 4 }x -3\]